### v2

@@ 3.2 Overlapping Strategy for Expert Parallelism
As mentioned in the previous section, the MoE operations in ScMoE architecture are completely decoupled from the backbone network, enabling parallel execution across two independent streams: one for the shared expert process and another for the MoE process. To enhance the efficiency, we implement asynchronous All-to-All communication operators to enable the overlapping of communication and computation within these streams, while computation operators are unable to execute concurrently due to the constraints on computing resources.
Figure 5: An overview of advanced expert parallelism using our ScMoE architecture and overlapping strategy. The red line represents the decoupled MoE stream and the numbers ① through ④ denote the potential locations for expert computation.
Figure 6: The timeline of different MoE architectures with corresponding parallel strategies, including pipeline and our proposed Overlapping. In each timeline, the length of each operator represents its time cost, while multiple rows indicate the utilization of parallel CUDA streams. The standard MoE utilizes top-2 gating, whereas the shared-expert MoE and ScMoE activate one shared expert alongside one gate-selected expert.
@@
$=\min_{\mathcal{K}}(|\sum_{i=1}^{\mathcal{K}-1}$ $COMP_{i}-\mathcal{T}_{disp}|+|\!\!\sum_{i=\mathcal{K}+1}^{4}\!\!\!COMP_{i}-\mathcal{T}_{comb}|)$
$\mathcal{T}_{overall}^{block}\geq|(\mathcal{T}_{comp}^{pre}+\mathcal{T}_{comp}^{post})-(\mathcal{T}_{disp}+\mathcal{T}_{comb})|$ (12)
$\mathcal{T}_{overall}^{block}\leq(\mathcal{T}_{comp}^{pre}+\mathcal{T}_{comp}^{post})+(\mathcal{T}_{disp}+\mathcal{T}_{comb})$ (13)
@@
The widely-used pipeline parallel strategy equally segments input tokens into smaller fine-grained chunks, enabling concurrent computation and communication dispatched on distinct GPU streams Hwang et al. (2022); Zhang et al. (2023). Contrary to standard MoE with pipelining (2nd timeline), our ScMoE with our proposed overlapping strategy (4th timeline) reduces the total communication time by half, leading to the same cost as the shared-expert MoE (3rd timeline). Our approach further optimizes performance by overlapping communication with the computation duration ($\mathcal{T}_{Atten}+\mathcal{T}_{SE}+\mathcal{T}_{MLP}$), which extends beyond the duration of expert computation achieved through pipelining.
@@
Our strategy possesses the capability to fully overlap communication if the communication tasks can be accommodated within the overlapping window. This advantage is not shared by the pipeline strategy as it cannot overlap initial and terminal data transmissions Huang et al. (2019); Narayanan et al. (2019). In cases where communication durations exceed the available overlap duration, our strategy can be augmented with pipelining (5th timeline), thereby utilizing the expert computation duration to further hide communication.
3.3 Expert Offloading Strategy for Memory-Limited Inference
Figure 7: Illustrations of various expert migration methods to improve the efficiency of expert offloading: (a) speculative expert migration, exemplified by Pre-gated MoE Hwang et al. (2023a), and (b) our ScMoE’s determinate expert migration. The red dashed line indicates expert selection and the transfer of expert parameters from CPU memory to GPU memory, while the black or red solid lines represent the data flow of representations processed by the Attention, MLP, and expert modules.
As discussed in Section 2.3, MoE models encounter significant deployment challenges during on-device inference due to their model size exceeding GPU memory capacity, necessitating the offloading of expert parameters to CPU memory. Moreover, decoder-only models use an autoregressive process for natural language generation (NLG) inference tasks, allowing for per-token processing of MoE. Specifically, only the two activated experts (top-2 gating) for each token need to be transferred from CPU to GPU memory for computation, thereby reducing peak GPU memory usage.
To minimize the overhead caused by migrating activated expert parameters from CPU to GPU, speculative expert migration methods, represented by Pre-gated MoE Hwang et al. (2023a), use preceding-layer representations to predict expert selection, as illustrated in Figure 7 (a). This approach allows the overlap of expert migration duration with the computation of preceding modules. Moreover, speculative expert migration methods adjust only the expert selection process, while expert computation continues along the same data flow of representations as in standard MoE.
In contrast, our proposed ScMoE architecture utilizes the gate-selected expert to compute the preceding-layer representations, inherently facilitating early expert migration well before the expert computation in the current layer. Based on this, we implement an expert offloading strategy that keeps non-expert and shared expert modules in GPU memory while offloading other gate-selected experts to CPU memory. After the Attention module in the preceding layer generates intermediate representations, the gate determines expert selection and issues asynchronous migration of the activated expert, as illustrated in Figure 7 (b). This approach allows expert migration to overlap with the computation duration ($\mathcal{T}_{Atten}+\mathcal{T}_{SE}+\mathcal{T}_{MLP}$). Importantly, expert selection in our method adheres to the logic of the pre-trained ScMoE model, without any speculation.
Additionally, existing expert migration methods cannot be adapted to overlap communication in expert parallelism. This is because they do not decouple dependencies in the data flow of expert processing representations, and therefore cannot adjust the All-to-All communication of these representations.
@@ 4.1 Experimental Setup
Hardware Configurations. To assess the effectiveness of our proposed overlapping strategy for enhancing expert parallelism, we conduct experiments on three hardware configurations: 8$\times$×A30-PCIe, 8$\times$×A800-NVLink and 16$\times$×A800-NVLink (across 2 nodes). These configurations cover scenarios with both high and low communication-to-computation ratios. Additionally, we evaluate our proposed expert offloading strategy on a configuration with a single A30-PCIe GPU.
Experiments on Vision Model. To evaluate the efficacy of our MoE architectures on vision tasks, we conduct experiments on SwinV2-MoE model, which is a state-of-the-art vision transformer model built upon the Tutel MoE framework Hwang et al. (2022); Liu et al. (2021). Specifically, we pre-train the SwinV2-MoE models with various MoE architectures on ImageNet-1K image classification dataset, and subsequently evaluate their accuracy on the corresponding test set. It is noteworthy that the integration of the MoE module within SwinV2 is confined to stages 3 and 4, with our architectural enhancements being selectively applied to the MoE modules in stage 3—the deepest submodel. Given our hardware constraints, we configure each MoE module with 8 experts, assigning one expert per GPU device.
Experiments on Language Model. For natural language generation (NLG) tasks, we utilize the standard implementations of GPT-2 Radford et al. (2019) and GPT-3 Brown et al. (2020) from Fairseq Ott et al. (2019), augmented with Tutel MoE to construct GPT2-MoE and GPT3-MoE models featuring an 8-expert MoE module. Specifically, we implement it by substituting the MLP with MoE in the second transformer block of every consecutive pair. Then, we pre-train the models with different architectures on the OpenWebtext dataset Gokaslan & Cohen (2019). In addition to assessing the pre-training validation perplexity, we also conduct a zero-shot evaluation of the pre-trained model on the WikiText-103 Merity et al. (2016) dataset.
More setup details are illustrated in Appendix A.6
Table 2: Test top-1 accuracy and end-to-end speedup of train and inference (one iteration) for SwinV2-MoE-S Hwang et al. (2022) models with various architectures pre-trained on ImageNet-1K for 90 epochs in 8$\times$×A30-PCIe scenario, using standard MoE with top-2 gating as the baseline.
@@
Our ScMoE 79.38% 1.43$\times$ 1.66$\times$
Table 3: Comparison of zero-shot perplexity on WikiText-103 and end-to-end speedup analysis of train and inference (one iteration) for our pre-trained GPT2-MoE-Medium Radford et al. (2019) models with various architectures in 8$\times$×A800-NVLink scenario, using standard top-2 MoE as the baseline.
Model WikiText-103 Train Inference
[…]
Standard top-2 MoE 19.18 1 1
Shared-Expert MoE 17.94 1.04$\times$ 1.06$\times$
Our ScMoE 17.62 1.12$\times$ 1.17$\times$
4.2 Analysis of Model Quality and Efficiency in Distributed Scenarios
In this section, we assess the model quality of our proposed ScMoE, which is a novel architectural design distinct from existing MoE models. Furthermore, we evaluate the efficiency of our ScMoE models in distributed scenarios, which are accelerated through our proposed overlapping strategy for enhancing expert parallelism. To maintain the same computational volume as the standard top-2 MoE, both the experimental Shared-Expert MoE and our ScMoE utilize one shared expert and one gate-selected expert.
@@ 4.2 Analysis of Model Quality and Efficiency in Distributed Scenarios
Figure 8: Overhead analysis for each pair of Block-MLP and Block-MoE within SwinV2-MoE-S model, deployed across three different distributed scenarios. “Topk” denotes the standard top-k MoE, while the one followed by the suffix “P” indicates using pipeline optimization as implemented by Tutel Hwang et al. (2022). “Top1+SE1” refers to the shared-expert MoE.
Refer to caption (a) GPT2-MoE-Small Refer to caption (b) GPT2-MoE-Medium
Figure 9: Token-wise validation perplexity curves for training MoE models with different MoE architectures.
4.2.1 Vision Model
Table 2 shows that our ScMoE and the standard top-2 MoE attain a comparable accuracy of 79.3%, while the shared-expert MoE delivers the highest accuracy, with a marginal increase of 0.2%. These methods, which utilize two activated experts, consistently outperform the standard top-1 MoE in terms of model quality, as the top-1 approach involves fewer parameters for computation.
Notably, in the 8$\times$×A30-PCIe scenario where communication overhead accounts for 60% of the total MoE time, our ScMoE model exhibits 30% speed improvement in training and 40% in inference compared to the standard top-2 MoE. Furthermore, it even surpasses the top-1 MoE by 13% in training speed and 20% in inference speed.
Consequently, our proposed ScMoE architecture improves the MoE models’ operational efficiency without compromising on model quality. Additional experiments with the SwinV2-MoE-B model demonstrate similar results, as detailed in Appendix A.4.
4.2.2 Language Model
As illustrated in Figure 9, both our ScMoE and the shared-expert MoE, which utilize one shared expert, demonstrate lower validation perplexity curves compared to the standard top-2 MoE during training on both the GPT2-MoE-Small and GPT2-MoE-Medium models. Consistently, in zero-shot evaluations on the WikiText-103 dataset (see Table 3), these two methods yield lower perplexity scores, with our ScMoE achieving 17.62 and the shared-expert MoE reaching 17.94, in contrast to the standard top-2 MoE, which results in a perplexity of 19.18.
As exemplified by the performance of GPT2-MoE-Small and GPT2-MoE-Medium models, our ScMoE performs high model quality across a range of network depths. Notably, our ScMoE demonstrates a superior training curve and zero-shot perplexity compared to the shared-expert MoE in the GPT2-MoE-Medium model, while the shared-expert MoE exhibits a slightly better training curve in the GPT2-MoE-Small model. This suggests that our ScMoE may offer greater benefits in deeper networks, aligning with our theoretical analysis of the shortcut connections, as detailed in Appendix A.1. Additional zero-shot results of GPT2-MoE-Small are presented in Appendix A.4.
In addition, our ScMoE achieves 11% speed improvement in training and 15% in inference compared to standard top-2 MoE in the 8$\times$×A800-NVLink scenario, where communication constitutes 15% of the total MoE time.
@@ 4.2.3 Analysis of Overhead and Acceleration
In addition to exhibiting the significant end-to-end speedup of our ScMoE in Tables 2 and 3, we delve into a detailed analysis of the overhead and the acceleration effect with our overlapping strategy, which can be generalized to other Transformer-based MoE models.
In the communication-intensive 8$\times$×A30-PCIe scenario (Figure 8a), our ScMoE overlaps 70% communication time, resulting in a 27% speed improvement over shared-expert MoE, a 42% improvement over the pipelined standard top-2 MoE, and a 15% improvement over the pipelined standard top-1 MoE. In the 8$\times$×A800-NVLink scenario (Figure 8b), which features almost minimal communication overhead, our approach maintains its acceleration by completely overlapping communication.
In the multi-node distributed scenario (Figure 8c), with 16$\times$×A800-NVLink across two nodes, communication incurs more significant overhead than in the single-node 8$\times$×A800-NVLink scenario due to the lower-bandwidth inter-node Ethernet Li et al. (2020). Here, our ScMoE achieves complete overlap, resulting in a 24% speed improvement over the shared-expert MoE, a 43% improvement over the pipelined standard top-2 MoE, and a 13% improvement over the pipelined standard top-1 MoE.
@@
4.2.4 Analysis of More Activated Experts
As increasing the number of activated experts within standard MoE is correlated with enhancements in model quality, we implement this augmentation in our ScMoE by increasing the count of gate-selected experts that process the preceding-layer representations, while maintaining the process of current-layer representations. To investigate the benefits of more activated experts, we implement the ScMoE-2, which employs top-2 experts for the preceding layer and one shared expert for the current layer.
Table 4: Comparison of validation perplexity and end-to-end speedup analysis of train and inference (one iteration) for our pre-trained GPT3-MoE-XL Brown et al. (2020) models with various architectures in 8$\times$×A800-NVLink scenario, using standard MoE with top-2 gating as the baseline. “ScMoE-2” refers to the activation of one shared expert and two gate-selected experts.
Model Validation Train Inference
[…]
Standard top-2 17.52 1 1
Our ScMoE 16.46 1.12$\times$ 1.18$\times$
[…]
Our ScMoE-2 16.27 1.05$\times$ 1.08$\times$
4.3 Analysis of Memory-Limited Inference
To tackle the challenges of on-device inference, where it is not feasible to store the entire MoE model parameters in GPU memory, our strategy involves offloading gate-selected experts to CPU memory while retaining non-expert and shared expert modules in GPU memory. As demonstrated in Table 10 (a), our expert offloading strategy reduces peak GPU memory usage by 50% for the GPT2-MoE-Medium model and by 60% for the GPT3-MoE-XL model when deployed in the inference scenario using a single A30-PCIe GPU. Furthermore, it is anticipated that models with more gate-selected experts in each MoE module will experience a larger percentage reduction in GPU memory usage.
Since the offloaded expert parameters must be loaded into the GPU memory for expert computation, the blocking execution of this expert migration results in significant overhead. As shown in Table 10 (b), the blocking expert migration introduces an additional overhead of 80% in GPT2-MoE-Medium and 240% in GPT3-MoE-XL, substantially increasing the MoE block latency. To mitigate this issue, our strategy of asynchronously executing the determinate expert migration effectively reduces the additional costs by 75% in GPT2-MoE-Medium and 25% in GPT3-MoE-XL.
Furthermore, it is evident that expanding the model size from Medium to XL significantly raises the cost proportion related to expert migration. This is because the per-token decoding process during inference is memory-bound Patel et al. (2024); Wu et al. (2024). The larger model size leads to a proportional increase in the duration of memory transfer, without a corresponding increase in computation time.
Figure 10: Peak GPU memory usage (a) and MoE block latency (b) for various memory-limited inference methods applied to the GPT2-MoE-Medium and GPT3-MoE-XL models using ScMoE. “GPU-only” indicates that the entire model is stored in GPU memory. “Offload” refers to our strategy of offloading expert parameters to CPU with blocking expert migration. “Offload-Async” denotes the use of asynchronous expert migration to overlap its duration.
Figure 11: Results from the analysis of the proposed shortcut connection, during the 90-epoch training (including a 10-epoch warm-up) of the SwinV2-MoE-S model Hwang et al. (2022). Employing the same MoE module to select the top-1 expert twice for processing each input token’s two representations from the current and preceding layers, respectively, (a) illustrates the percentage of tokens that retain the same expert selection across the current layer and preceding layer, (b) shows the L2 distance between these two representations. Using the DGMoE, which imposes a constraint against repeatedly selecting the same expert, (c) presents the average gating score for the preceding-layer representations, (d) displays the average gating score for the current-layer representations.
4.4 Analysis of the Proposed Shortcut Connection
Our comprehensive empirical results have demonstrated that our ScMoE architecture facilitates efficiency optimizations without compromising model quality. We further explore the proposed shortcut connection in greater depth, uncovering the reasons behind its algorithmic effectiveness.

### v3

@@ 3.2 Overlapping Strategy for Expert Parallelism
As mentioned in the previous section, the MoE operations in ScMoE architecture are completely decoupled from the backbone network, enabling parallel execution across two independent streams: one for the shared expert process and the other for the MoE process. To enhance efficiency, we implement asynchronous All-to-All communication operators to enable the overlapping of communication and computation within these streams, while computation operators are unable to execute concurrently due to the constraints on computing resources.
@@
$=\min_{\mathcal{K}}(|\sum_{i=1}^{\mathcal{K}-1}$ $COMP_{i}-\mathcal{T}_{disp}|+|\!\!\sum_{i=\mathcal{K}+1}^{4}\!\!\!COMP_{i}-\mathcal{T}_{comb}|),$
$\mathcal{T}_{overall}^{block}\geq|(\mathcal{T}_{comp}^{pre}+\mathcal{T}_{comp}^{post})-(\mathcal{T}_{disp}+\mathcal{T}_{comb})|,$ (12)
$\mathcal{T}_{overall}^{block}\leq(\mathcal{T}_{comp}^{pre}+\mathcal{T}_{comp}^{post})+(\mathcal{T}_{disp}+\mathcal{T}_{comb}).$ (13)
@@
Figure 6: An overview of advanced expert parallelism using our proposed ScMoE architecture and overlapping strategy. The red line represents the decoupled MoE stream and the numbers ① through ④ denote the potential locations for expert computation.
Figure 7: The timeline of different MoE architectures with corresponding parallel strategies, including pipeline and our proposed overlapping strategy. In each timeline, the length of each operator represents its time cost, while multiple rows indicate the utilization of parallel CUDA streams. The standard MoE utilizes top-2 gating, whereas the shared-expert MoE and ScMoE activate one shared expert alongside one gate-routed expert.
@@
The widely-used pipeline parallel strategy equally segments input tokens into smaller fine-grained chunks, enabling concurrent computation and communication dispatched on distinct GPU streams (Hwang et al., 2023; Zhang et al., 2023). Contrary to standard MoE with pipelining (2nd timeline), our proposed ScMoE with the overlapping strategy (4th timeline) significantly reduces the total time. This reduction is attributed to the decrease in absolute communication time, similar to that in the shared-expert MoE (3rd timeline), and the overlap of communication with the computation duration ($\mathcal{T}{Atten}+\mathcal{T}{SE}+\mathcal{T}_{MLP}$), which extends beyond the overlap duration achieved through pipelining.
@@
Our strategy possesses the capability to fully overlap communication if the communication can be accommodated within the overlapping window. This advantage is not shared by the pipeline strategy as it cannot overlap the initial and terminal data transmissions (Huang et al., 2019; Narayanan et al., 2019). In cases where communication durations exceed the available overlap duration, our strategy can be augmented with pipelining (5th timeline), thus utilizing the expert computation duration to further hide communication.
@@ 4.1 Experimental Setup
To assess the effectiveness of our proposed overlapping strategy for enhancing expert parallelism, we conduct experiments on three hardware configurations: 8$\times$×A30-PCIe, 8$\times$×A800-NVLink, and 16$\times$×A800-NVLink (across 2 nodes). These configurations cover scenarios with both high and low communication-to-computation ratios. Furthermore, we evaluated our methods using both vision models (SwinV2-MoE) (Hwang et al., 2023) and language models (GPT2-MoE, GPT3-MoE, LLaMA2-MoE) (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023). Additional details on the experimental setup are provided in Appendix A.8.
4.2 Analysis of Model Quality and Efficiency
In this section, we assess the quality of the models with our proposed ScMoE architecture. Furthermore, we evaluate the efficiency of ScMoE models in distributed scenarios, which are accelerated through our proposed overlapping strategy for enhancing expert parallelism. To maintain the same computational volume as the standard top-2 MoE, both the experimental shared-expert MoE and our ScMoE utilize one shared expert and one gate-routed expert.
Table 1: Test top-1 accuracy and end-to-end speedup of train and inference (one iteration) for SwinV2-MoE-S (Hwang et al., 2023) models with various architectures pre-trained on ImageNet-1K for 90 epochs in the 8$\times$×A30-PCIe scenario, using standard MoE with top-2 gating as the baseline.
@@ 4.2 Analysis of Model Quality and Efficiency
Our ScMoE (Pos-1) 79.14% 1.36$\times$ 1.54$\times$
Our ScMoE (Pos-2) 79.38% 1.43$\times$ 1.66$\times$
Our ScMoE (Pos-3) 79.20% 1.49$\times$ 1.82$\times$
4.2.1 Vision Model
Table 1 shows that ScMoE (Pos-2) and the standard top-2 MoE attain a comparable accuracy of 79.3%, while the shared-expert MoE delivers the highest accuracy, with a marginal increase of 0.2%. In 8$\times$×A30-PCIe where communication overhead accounts for 60% of the total MoE time, ScMoE (Pos-2) exhibits 30% speed improvement in training and 40% in inference compared to the standard top-2 MoE.
In scenarios characterized by a high communication-to-computation ratio, where communication is hard to be completely overlapped by computation, ScMoE architectures with extended overlap durations can achieve superior speedup. Specifically, ScMoE (Pos-3), which has the longest overlap duration of $\mathcal{T}_{Atten}+\mathcal{T}_{SE}+\mathcal{T}_{MLP}$, achieves the highest acceleration, with a 1.49× speedup in training and a 1.82× speedup in inference. Furthermore, the three different ScMoE architectures result in minimal variations in accuracy, ranging from 79.14% to 79.38%. Additionally, these methods, which utilize two activated experts, consistently outperform the standard top-1 MoE in terms of model quality, as the top-1 approach activates fewer parameters.
4.2.2 Language Model
To demonstrate the effectiveness of our proposed ScMoE architecture in models with two prevailing MoE placement frequency designs, we perform experiments using the GPT2-MoE-Medium and LLaMA2-MoE models, positioning the MoE module in every second Transformer block for GPT2-MoE and in every Transformer block for LLaMA2-MoE. Specifically, we utilize the ScMoE architecture with the configuration of “Pos-2” on the GPT2-MoE model, since this setup yields the lowest final validation loss in our experiments across various shortcut-connected positions, as discussed in Section 5.2. Furthermore, the ScMoE in LLaMA2-MoE adopts the “Pos-1” configuration, as this setup has already maximized the potential overlap duration in the scenario of MoE placement in every block.
As shown in Table 2, our ScMoE models achieve the highest average scores, with 38.69 in GPT2-MoE and 38.96 in LLaMA2-MoE. Furthermore, when integrated with our ScMoE, GPT2-MoE experienced an 11% improvement in training speed and an 15% improvement in inference speed compared to the standard top-2 MoE, in the 8$\times$×A800-NVLink scenario where communication accounts for 15% of the total MoE time. In LLaMA2-MoE, our ScMoE accelerated training by 1.14$\times$ and inference by 1.21$\times$, demonstrating superior efficiency compared to other methods.
Table 2: Comparison of zero-shot evaluation and end-to-end speedup of training and inference (one iteration) for the pre-trained GPT2-MoE and LLaMA2-MoE models with various architectures in the 8$\times$×A800-NVLink scenario, using standard top-2 MoE as the baseline.
Model Method Train Inference HellaSwag PIQA WinoGrande BoolQ ARC-E OBQA RACE MathQA AVG.($\uparrow$)
GPT2-MoE Standard top-2 1 1 27.53 59.19 48.62 59.72 38.43 25.20 23.83 20.37 37.86
Shared-Expert 1.04$\times$ 1.06$\times$ 27.23 59.09 51.22 60.00 38.85 26.60 25.07 20.57 38.58
Our ScMoE 1.12$\times$ 1.17$\times$ 27.70 59.25 52.09 60.76 39.23 25.40 24.98 20.10 38.69
LLaMA2-MoE Standard top-2 1 1 28.40 60.07 50.83 58.26 38.72 24.60 25.17 21.07 38.39
Shared-Expert 1.06$\times$ 1.11$\times$ 29.08 60.01 50.91 60.92 38.59 24.80 25.45 20.80 38.82
Our ScMoE 1.14$\times$ 1.21$\times$ 29.09 60.55 51.38 57.25 38.89 26.40 27.08 21.07 38.96
@@ 4.2.2 Language Model
Figure 8: Overhead analysis for each pair of Block-MLP and Block-MoE within SwinV2-MoE-S model, deployed across three different distributed scenarios. “Topk” denotes the standard top-k MoE, while the one followed by the suffix “P” indicates using pipeline optimization as implemented by Tutel (Hwang et al., 2023). “Top1+SE1” refers to the shared-expert MoE.
Figure 9: Results from the analysis of the proposed shortcut connection, during the 90-epoch training (including a 10-epoch warm-up) of the SwinV2-MoE-S model (Hwang et al., 2023). Employing the same MoE module to select the top-1 expert twice for processing each input token’s two representations from the current and preceding layers, respectively, (a) illustrates the percentage of tokens that retain the same expert selection across the current layer and preceding layer, (b) shows the L2 distance between these two representations. Using the DGMoE, which imposes a constraint against repeatedly selecting the same expert, (c) presents the average gating score for the preceding-layer representations, (d) displays the average gating score for the current-layer representations.
@@ 4.2.3 Analysis of Overhead and Acceleration
In addition to exhibiting the end-to-end speedup of ScMoE in Tables 1 and 2, we delve into a detailed analysis of the overhead and the acceleration effect with our overlapping strategy, which can be generalized to other MoE models.
In the communication-intensive 8$\times$×A30-PCIe scenario (Figure 8(a)), our ScMoE overlaps 70% communication time, resulting in a 27% speed improvement over shared-expert MoE, a 42% improvement over the pipelined standard top-2 MoE, and a 15% improvement over the pipelined standard top-1 MoE. In the 8$\times$×A800-NVLink scenario (Figure 8(b)), which features almost minimal communication overhead, our approach maintains its acceleration by fully overlapping.
In multi-node scenario (Figure 8(c)), with 16$\times$×A800-NVLink across two nodes, communication incurs more significant overhead than in the single-node 8$\times$×A800-NVLink scenario due to the lower-bandwidth inter-node Ethernet (Li et al., 2020). Here, our ScMoE achieves complete overlap, resulting in a 24% speed improvement over the shared-expert MoE, a 43% improvement over the pipelined standard top-2.
@@
5 Discussion
The empirical results presented in Section 4.2 have demonstrated that our proposed ScMoE architecture facilitates efficiency optimizations without compromising model quality. Subsequently, we delve into a more thorough examination of the proposed shortcut connection, uncovering potential underlying reasons for its algorithmic effectiveness, and identifying opportunities for further development.
5.1 Delve into the Proposed Shortcut Connection
5.1.1 Analysis of Gating Behaviors